OpenStudy (anonymous):

12 - |n + 4| = 20 is no solution??

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

12 - |n+4| = 20 -|n+4| = 8 n + 4 = -8 n = -12 n + 4 = 8 n = 4

OpenStudy (anonymous):

what about after isolating the absolute?

OpenStudy (anonymous):